# Unsupervised Episode Detection for Large-Scale News Events

### v1

@@ Abstract.
Episodic structures are inherently interpretable and adaptable to evolving large-scale key events. However, state-of-the-art automatic event detection methods overlook event episodes and, therefore, struggle with these crucial characteristics. This paper introduces a novel task, episode detection, aimed at identifying episodes from a news corpus containing key event articles. An episode describes a cohesive cluster of core entities (e.g., “protesters”, “police”) performing actions at a specific time and location. Furthermore, an episode is a significant part of a larger group of episodes under a particular key event. Automatically detecting episodes is challenging because, unlike key events and atomic actions, we cannot rely on explicit mentions of times and locations to distinguish between episodes or use semantic similarity to merge inconsistent episode co-references. To address these challenges, we introduce EpiMine, an unsupervised episode detection framework that (1) automatically identifies the most salient, key-event-relevant terms and segments, (2) determines candidate episodes in an article based on natural episodic partitions estimated through shifts in discriminative term combinations, and (3) refines and forms final episode clusters using large language model-based reasoning on the candidate episodes. We construct three diverse, real-world event datasets annotated at the episode level. EpiMine outperforms all baselines on these datasets by an average 59.2% increase across all metrics.
Unsupervised Episode Detection, Event Detection, Segment Classification, Large Language Models, Key Term Extraction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Information extractionccs: Information systems Data mining
Figure 1. Example of an event structure hierarchy. We take a key event corpus as input and aim to detect each episode node with a corresponding cluster of relevant text segments.
1. Introduction
@@
Given the saturation of real-time news accessible at our fingertips, reading and processing the critical information of a key event has become an increasingly daunting challenge. Consequently, research on automatic textual event detection has recently attempted to integrate the manner in which humans neurologically perceive/store events into textual event detection methods. Specifically, neuroscientists studying event representations in human memory find that events are stored in a top-to-bottom hierarchy, as demonstrated in Figure 1. The deeper the hierarchical event level, the more fine-grained its corresponding text granularity (Zhang et al., 2022): we consider a theme as corpus-level (all articles discussing the 2019 Hong Kong Protests), key event as document-level (an article typically discusses a full one to two day key event), episode as segment-level, and atomic action as sentence or phrase-level.
@@
Furthermore, neurological research (Baldassano et al., 2017; Khemlani et al., 2015) indicates that events are encoded into memory as episodic structures. Representing events as discrete episodes helps us piece together a coherent narrative by considering the sequence of actions, reactions, and developments over time. This empowers several downstream tasks (e.g., event schema generation, event prediction), which may benefit from insights into the causes and consequences of events that might be missed with a more generalized analysis. Despite its strengths, existing automatic event extraction works fail to consider the episode-level.
@@
For example, key event detection specifically seeks to output “a set of thematically coherent documents” for each key event (Zhang et al., 2022; Liu et al., 2023). However, it is challenging to manually parse through a large cluster of relevant articles in order to gain an efficient and compact understanding of a given event. To address the lack of interpretability of such article clusters, the adjacent task of timeline summarization (Steen and Markert, 2019; Li et al., 2021b; Gholipour Ghalandari and Ifrim, 2020; Chen et al., 2023) aims to identify the dates and a compact summary for each key event. However, high-level timelines are typically applicable for historical themes and thus unrealistic for currently evolving key events where fine-grained timeline summarization is more suitable. Hence, event chain mining (Jiao et al., 2023) attempts to address this by mining a series of temporally-ordered atomic actions at the phrase-level; however, the granularity level is often too fine-grained to properly represent a large-scale key event. Thus, we aim to tackle the novel task of episode detection to pave the way for a more effective event representation.
@@
Episode detection aims to detect episodes from a news corpus containing key event articles. An episode can be described as a cohesive cluster of subjects performing actions at a certain time and location, occurring as part of a larger sequence of episodes under a specific key event. Episode detection introduces a unique set of challenges, which we address using our novel framework, EpiMine. EpiMine is an unsupervised episode detection framework that automatically detects meaningful episodic events and their corresponding text segments in a large key event corpus, all without any level of human supervision or labeled training data. EpiMine is comprised of the following components: (1) discriminative co-occurrence detection, (2) episode partitioning, (3) candidate episode estimation, and (4) episode-segment classification. Collectively, they tackle the unique challenges of episode detection, detailed below:
@@
Challenge 1: Journalists do not timestamp episodes. Key event detection partitions a thematic corpus into document-level clusters by heavily relying on explicit temporal features, like publication dates (Zhang et al., 2022). For example, an article primarily discusses one key event, which can then be roughly mapped to the article’s publication date. However, this assumption fails at the episode-level, where there is no guarantee to have a distinct timestamp associated with each text segment that discusses a new episode. Fortunately, we can take advantage of the idea that journalists naturally partition news articles by sequentially discussing distinct episodes: {adjustwidth}1em0em Example: An article likely completes its discussion of the episode A, protesters storming the Legislaive Council, before episode B, “protesters vandalized the Legislative Chamber” (Figure 3). Hence, in order to partition articles into distinct episode segments, EpiMine must identify whether or not two consecutive segments are discussing the same or different episodes, which brings us to our next challenge.
@@
Challenge 2: Episodes contain semantically diverse actions. Each episode features a set of unique atomic actions, which we can utilize for determining whether or not two segments discuss the same episode. However, for clustering actions, existing methods (Jiao et al., 2023) rely heavily on semantic similarity. This not realistic for episode-segment clustering: {adjustwidth}1em0em Example:protesters spray-painted slogans” and “they unfurled the colonial-era flag” will fall under the same episode, but are semantically different and unlikely to be clustered. Alternatively, we can identify salient terms that reflect the same episode (“barriers” and “shoved” are unique to Episode A; “defaced” and “walls” for Episode B), by exploiting corpus-level signals. Specifically, if we frequently see “defaced” mentioned together with “walls” (or their respective synonyms) and not with other terms (and vice-versa), then we consider them a discriminative co-occurrence. Consequently, when such co-occurrences remain consistent across text segments, this indicates the same episode being discussed. Conversely, if a sufficient shift in the combination of terms occurs, then this indicates that a different episode is being discussed.
@@
Challenge 3: Articles often do not feature all episodes. Given the real-time nature of reporting, an article may feature only a subset or none of ground-truth episodes from a multi-day key event (e.g., focusing on a high-level analysis of the key event or mentioning other related key events). To minimize this noise, EpiMine seeks to identify the set of articles which maximizes the quantity and quality of potential episodes. It then merges any article partitions across these articles which likely discuss the same episode and employs a large language model (LLM) to provide a more fluent interpretation of the candidate episodes, accounting for the episode’s core entity, actions, object, location, and time period. This allows EpiMine to finally map the remaining non-salient article segments to these episodes, pruning any candidates which are not sufficiently supported by the remaining articles. We summarize our core contributions:
  1. (1)
@@
    We introduce the novel task of episode detection, which takes in a key-event corpus and outputs multiple detected episodes and their related segments extracted from the corpus.
  2. (2)
    We propose a novel unsupervised episode detection method, EpiMine, which exploits discriminative term co-occurrences to estimate candidate episode partitions from top articles. It refines these candidate episodes using LLM-based reasoning to output a comprehensive and diverse set of episodes.
  3. (3)
@@
    We construct three novel datasets, reflecting a diverse set of real-world themes and thirty global key events, as no large-scale key event-specific news corpus exists for this task where the key events are guaranteed to contain distinguishable episodes.
  4. (4)
    We compare EpiMine with five other baselines through an extensive set of experiments and case studies performed on our real-world datasets, demonstrating that it outperforms all baselines by, on average, a 59.2% increase across all metrics.
@@
2. Related Works
2.1. Event Extraction
Understanding events in unstructured texts is crucial, leading to substantial research in event detection and extraction (Liu et al., 2018a; Du and Cardie, 2020; Li et al., 2021a; Lu et al., 2021; Qi et al., 2022; Jiao et al., 2022). Existing work, such as Han et al. (2019); Wang et al. (2020); Ahmad et al. (2021), have focused on event relation extraction and prediction, while salient event identification remains a significant area of study (Liu et al., 2018b; Jindal et al., 2020; Wilmot and Keller, 2021). Furthermore, event process understanding has garnered recent attention (Zhang et al., 2020; Chen et al., 2020). However, these studies typically require costly expert annotations. Some efforts have sought to mitigate this challenge through unsupervised approaches (Weber et al., 2018; Li et al., 2020). The pioneering work (Chambers and Jurafsky, 2008) introduced event chains as a novel representation of structured news knowledge. In the realm of biological reading comprehension, Berant et al. (2014) extracted events and their relationships in biological processes. More recently, Zhang et al. (2021) utilized event graphs constructed from large corpora to derive event chains for narrative understanding. Despite these advancements, most existing studies concentrate on phrase-level or sentence-level events (analogous to actions in this paper), neglecting the hierarchical structures of events at varying granularities within textual data. Additionally, these approaches often rely on human-curated event ontologies, are restricted to predefined event types, and fail to capture events from open-domain texts. Although key event extraction has been explored in some studies (Zhang et al., 2022; Jiao et al., 2023), the automatic detection of episodes remains under-explored and critically important.
2.2. Timeline Summarization
Timeline summarization (TLS) automatically identifies key dates of major events and provide concise descriptions of occurrences on those dates. Previous approaches to TLS have predominantly focused on extractive methods. For instance, Nguyen et al. (2014) built thematic timelines for general-domain topics by selecting and ranking events relevant to an input query. Martschat and Markert (2018) introduced a temporally sensitive submodularity framework, designing objective functions and constraints that model the temporal dimension inherent in timeline summarization. Steen and Markert (2019) were the first to propose an abstractive timeline summarization system. Their approach clusters sentences likely describing the same event and uses multi-sentence compression to generate summary sentences for each cluster. More recently, Li et al. (2021b) represented news articles as event-graphs, transforming the summarization task into compressing the graph to its salient sub-graph. Additionally, You et al. (2022) presented a joint learning-based heterogeneous graph attention network, integrating event detection into a unified framework to eliminate redundant sentences. While these methods involve key event grouping and understanding, they typically apply to historical themes, producing higher-level timelines. However, these approaches may not be practical for ongoing news events, where a detailed timeline summary (distinguishing between episodes) is more appropriate.
2.3. Topic Discovery
The goal of topic discovery is to group news articles by their themes, an area of study dating back to the early days of automated event theme identification from textual data (Allan et al., 1998). One approach (Zhou et al., 2015; Beykikhoshk et al., 2018) treats this task as a topic modeling issue, using LDA (Blei et al., 2003) or its enhanced versions. Another line of research (Sayyadi et al., 2009; Sayyadi and Raschid, 2013) constructs a graph based on keyword frequency and identifies themes from keyword clusters, mapping these themes to documents via keyword-driven feature similarities. However, these techniques do not consider temporal information, limiting them to discerning different thematic events. To address this, some studies (Fung et al., 2005; He et al., 2007) introduce tracking keyword spikes to pinpoint events. Specifically, (Zhao et al., 2012) uses burst detection (Kleinberg, 2004) to identify and utilize bursty terms for document representation and clustering. Building on this, (Ge et al., 2016) creates a burst information network with recognized bursts and their attributes for clustering. Finally, (Yoon et al., 2023) proposes unsupervised stream-based story discovery, which computes article embeddings their shared temporal themes. Overall, these methods solely focus on the key-event or atomic action level.
Figure 2. We detail the overall framework of EpiMine.
3. Preliminaries
3.1. Problem Definition
Definition 3.1 (Episode).
@@ 3.1. Problem Definition
An episode $E_{i}$ is one of a partially ordered sequence of subevents, $\{E_{1},\ldots,E_{i},\ldots E_{k}\}$, of a key (major) event $E$, where typically $2\geq k\geq 20$, and $E_{i}$ does not overlap with $E_{j}$ if $i\neq j$. Actions in the episode $E_{i}$ can be either semantically similar or diverse, but typically have relatively tight time, location, and/or thematic ( entities, actions, objects) proximity between them.
@@
EpiMine aims to extract episodes from a news corpus, where an episode occurs as a significant component of a larger group of episodes that fall under a specific key event. For instance, in Figure 1, without knowing Episode #1, “Protesters stormed the Legislative Council Complex”, readers would not have fully understood Episode #3, “Police dispersed protesters”. Hence, episodes help us understand the overall key event and are especially useful for events that are currently evolving, where finer context is required for sufficiently understanding them.
Definition 3.2 (Episode Detection).
@@
Given a corpus $\mathcal{D}$ about one key event, where each document $d\in\mathcal{D}$ is a news article, the goal of the task is to obtain the set of text segment clusters =$\mathcal{E}=\{E_{1},E_{2},\ldots,E_{K}\}$. Each episode cluster $E_{i}\subset\mathcal{S}=\{s_{1}^{1},s_{2}^{1},\ldots,s_{|d|}^{|\mathcal{D}|}\}$, where $\mathcal{S}$ contains all of the text segments identified in each document $d\in\mathcal{D}$, and every two clusters do not have overlapping text segments (i.e., $E_{i}\cap E_{j}=\emptyset$ for $i\neq j$).
@@
It is important to note that the number of episodes $K$ is not known in advance and oftentimes, a news article segment may discuss either episodes of a different key event (e.g., an episode with similar aspects that occurred in a different historical key event) or multiple episodes of the current key event. Nonetheless, our goal is to detect the most relevant episodes to the current key event at hand and consequently mine the most distinctive text segments for each of these (hence our constraint of $E_{i}\cap E_{j}=\emptyset$ for $i\neq j$).
3.2. Key Event News Article Pre-Processing
Given that the expected output for the episode detection task is a cluster of text segments, we first must segment each key event news article. We would like to ideally preserve both the primary aspects (e.g., core entities and their actions) and peripheral aspects (e.g., reactions to a core entity’s action) relevant to that episode, which may be helpful for cross-document episode co-reference resolution. In order to do this, we utilize the text segmentation method, C99 (Choi, 2000). Furthermore, in order to assist with the cohesiveness of the segment, we employ entity co-reference resolution before performing segmentation, which assists with retaining the context across text segments (“They surrounded the legislative building […]” $\rightarrow$ “The protesters surrounded the legislative building […]”). In the following section, we present our core methodology given these text segments (in their raw form, without co-references resolved) and their source articles as the primary inputs.
4. Methodology

### v2

# Synergizing Unsupervised Episode Detection with LLMs for Large-Scale News Events

@@ Abstract
State-of-the-art automatic event detection struggles with interpretability and adaptability to evolving large-scale key events—unlike episodic structures, which excel in these areas. Often overlooked, episodes represent cohesive clusters of core entities performing actions at a specific time and location; a partially ordered sequence of episodes can represent a key event. This paper introduces a novel task, episode detection, which identifies episodes within a news corpus of key event articles. Detecting episodes poses unique challenges, as they lack explicit temporal or locational markers and cannot be merged using semantic similarity alone. While large language models (LLMs) can aid with these reasoning difficulties, they suffer with long contexts typical of news corpora. To address these challenges, we introduce EpiMine, an unsupervised framework that identifies a key event’s candidate episodes by leveraging natural episodic partitions in articles, estimated through shifts in discriminative term combinations. These candidate episodes are more cohesive and representative of true episodes, synergizing with LLMs to better interpret and refine them into final episodes. We apply EpiMine to our three diverse, real-world event datasets annotated at the episode level, where it achieves a 59.2% average gain across all metrics compared to baselines.
Synergizing Unsupervised Episode Detection with LLMs
for Large-Scale News Events
Priyanka Kargupta  Yunyi Zhang  Yizhu Jiao  Siru Ouyang  Jiawei Han  Siebel School of Computing and Data Science  University of Illinois at Urbana-Champaign {pk36,yzhan238,yizhuj2,siruo2,hanj}@illinois.edu
1 Introduction
@@
Given the saturation of real-time news accessible at our fingertips, reading and processing a key event’s critical information has become an increasingly daunting challenge. Consequently, recent work on automatic textual event detection has attempted to integrate the manner in which humans neurologically perceive/store events into textual event detection methods. Specifically, neuroscientists studying event representations in human memory find that events are stored in a top-to-bottom hierarchy, as demonstrated in Figure 1. The deeper the hierarchical event level, the more fine-grained its corresponding text granularity Zhang et al. (2022): we consider a theme as corpus-level (all articles discussing the 2019 Hong Kong Protests), key event as document-level (an article typically discusses a full one to two day key event), episode as segment-level, and atomic action as sentence or phrase-level.
Figure 1: Example event structure hierarchy. Given a key event node’s corpus, detect its episode children and their respective relevant text segments.
@@
Furthermore, neurological research Baldassano et al. (2017); Khemlani et al. (2015) indicates that events are encoded into memory as episodic structures. Representing events as discrete episodes helps us piece together a coherent and concise narrative by focusing on meaningful clusters of actions, reactions, and developments, rather than examining each in isolation or as a whole. Despite its strengths, existing automatic event extraction works fail to consider the episode-level.
@@
For instance, key event detection focuses on identifying “a set of thematically coherent documents” for each key event Zhang et al. (2022); Liu et al. (2023), but manually parsing large clusters of articles is inefficient and lacks interpretability. Timeline summarization Steen and Markert (2019); Li et al. (2021a); Gholipour Ghalandari and Ifrim (2020); Chen et al. (2023) addresses this by providing dates and compact summaries, yet it suits historical themes better than evolving key events that require finer granularity. Event chain mining Jiao et al. (2023) takes a fine-granularity approach by identifying temporally ordered atomic actions, but its phrase-level granularity is often too fine and practically redundant for large-scale events (e.g., in Figure 1, the actions all describe the same episode). To bridge this gap, we propose the novel task of episode detection to pave the way for a more effective event representation.
@@
Episode detection aims to detect episodes from a news corpus containing key event articles. An episode can be described as a cohesive cluster of potentially diverse subjects performing actions at a certain time and location, occurring as part of a larger sequence of episodes under a specific key event. We introduce EpiMine, which detects meaningful episodic events and their corresponding text segments in a large key event corpus, all without any level of human supervision or labeled training data. EpiMine consists of: (1) episode indicative term mining, (2) episode partitioning, (3) LLM-enhanced episode estimation, and (4) episode-segment classification. Collectively, they tackle the unique challenges of episode detection, detailed below:
@@
Challenge 1: Episodes are not timestamped. Key event detection partitions a thematic corpus into document-level clusters by heavily relying on explicit temporal features, like publication dates, being associated with the key event articles Zhang et al. (2022). However, this assumption fails at the episode-level, where there is no guarantee to have a distinct timestamp associated with each text segment that discusses a new episode. Fortunately, we can take advantage of the idea that journalists naturally partition news articles by sequentially discussing distinct episodes: {adjustwidth}0.5em0em Example: An article likely completes its discussion of the episode A, protesters storming the Legislative Council, before episode B, “protesters vandalized the Legislative Chamber” (Figure 3). Hence, to partition articles into distinct episode segments, EpiMine must identify whether two consecutive segments are discussing the same or different episodes— bringing us to our next challenge.
@@
Challenge 2: Episodes contain semantically diverse actions. Each episode features a set of unique atomic actions, which can help determine if two segments discuss the same episode. However, for clustering actions, existing methods Jiao et al. (2023) rely heavily on semantic similarity. This is not realistic for episode-segment clustering: {adjustwidth}0.5em0em Example:protesters spray-painted slogans” and “they unfurled the colonial-era flag” will fall under the same episode, but are semantically different and unlikely to be clustered. Alternatively, we can identify notable (salient) terms unique to the same episode (episode A: “barriers” and “shoved”; episode B: “defaced” and “walls”), by exploiting corpus-level signals. For example, if “defaced” and “walls” are frequently mentioned together across the corpus (or their respective synonyms) and not with other terms, then they are a discriminative co-occurrence. When terms between two segments discriminatively co-occur, this indicates the same episode is being discussed. Conversely, if a sufficient shift in term combinations occurs, then a different episode is being discussed.
@@
Challenge 3: Articles often do not feature all episodes. Real-time news reporting often provides an incomplete coverage of multi-day events, with individual articles potentially omitting or partially addressing key episodes. Consequently, while LLMs could assist with the first two challenges, requiring multiple articles hinders their use given their long context limitations Li et al. (2024); Liu et al. (2024). To address this challenge, EpiMine seeks to select a minimal set of articles that maximizes both the quantity and quality of event episodes. It then merges any article partitions across these articles which likely discuss the same episode and synergizes with an LLM to provide a more fluent interpretation of the candidate episodes, accounting for the episode’s core entity, actions, object, location, and time period. This allows EpiMine to finally map the remaining non-salient article segments to these episodes, pruning any candidates which are not sufficiently supported by the remaining articles. We summarize our core contributions:
@@
  • Episode detection: novel task to detect episodes & their segments from a key-event corpus.
  • EpiMine, an unsupervised episode detection method which introduces discriminative term co-occurrence and episode partitioning.
@@
  • Three novel datasets, reflecting a diverse set of real-world themes and thirty global key events (no key event corpus exists for this task).
  • EpiMine outperforms all baselines by, on average, a 59.2% increase across all metrics.
@@
2 Related Works
2.1 Event Extraction
Event extraction has been widely studied, focusing on event detection Liu et al. (2018a); Du and Cardie (2020); Li et al. (2021b); Lu et al. (2021); Qi et al. (2022); Jiao et al. (2022), event relation extraction Han et al. (2019); Wang et al. (2020); Ahmad et al. (2021), and salient event identification Liu et al. (2018b); Jindal et al. (2020); Wilmot and Keller (2021). Recent work has also addressed event process understanding Zhang et al. (2020); Chen et al. (2020), though these often rely on expensive expert annotations. Some studies have introduced unsupervised methods to address annotation challenges Weber et al. (2018); Li et al. (2020). Some overlapping work exists in topic discovery, where Yoon et al. (2023) proposes unsupervised stream-based story discovery— computing article embeddings based on their shared temporal themes. Recently, large language models have demonstrated powerful general and event extraction-specific reasoning abilities Pai et al. (2024); Gao et al. (2024).
However, traditional and LLM-driven methods either, (1) focus on phrase/sentence-level events (analogous to actions in Figure 1), or (2) require human-curated event ontologies, often overlooking interpretable, yet meaningful granularities and open-domain texts, which go beyond pre-defined event types. While unsupervised granular event extraction has been explored Zhang et al. (2022); Jiao et al. (2023) at the document and phrase-level, episode detection is a more interpretable granularity that remains a largely unexplored, yet vital area.
2.2 Timeline Summarization
Timeline summarization (TLS) identifies key dates and concise descriptions for major events. Early methods were extractive, focusing on ranking events for thematic timelines Nguyen et al. (2014) or using submodular frameworks to model temporal dimensions Martschat and Markert (2018). Abstractive methods later emerged, such as sentence clustering and multi-sentence compression Steen and Markert (2019). More recent approaches are graph-based, such as event-graph representations for salient sub-graph compression Li et al. (2021a) and heterogeneous GATs for redundancy reduction You et al. (2022). While they effectively summarize key events as high-level timelines, they focus on historical themes. Episode-level timelines for ongoing news remain underexplored.
Figure 2: We detail the overall framework of EpiMine.
3 Preliminaries
3.1 Problem Definition
Definition 1 (Episode).
@@ 3.1 Problem Definition
An episode $E_{i}$ is one of a partially ordered sequence of subevents, $\{E_{1},\ldots,E_{i},\ldots E_{k}\}$, of a key (major) event $E$, where typically $2\leq k\leq 20$, and $E_{i}$ does not overlap with $E_{j}$ if $i\neq j$. Actions in the episode $E_{i}$ can be either semantically similar or diverse, but typically have relatively tight time, location, and/or thematic (entities, actions, objects) proximity.
@@
EpiMine aims to extract episodes from a news corpus, where an episode occurs as a significant component of a larger group of episodes that fall under a specific key event. For instance, in Figure 1, without knowing Episode #1, “Protesters stormed the Legislative Council Complex”, readers would not fully understand Episode #3, “Police dispersed protesters”. Hence, episodes help us understand the overall key event and are especially useful for events that are currently evolving, where finer context is required for sufficiently understanding them.
Definition 2 (Episode Detection).
@@
Given a corpus $\mathcal{D}$ about one key event, where each document $d\in\mathcal{D}$ is a news article, the task is to obtain a set of text segment clusters =$\mathcal{E}=\{E_{1},E_{2},\ldots,E_{k}\}$. Each episode cluster $E_{i}\subset\mathcal{S}=\{s_{1}^{1},s_{2}^{1},\ldots,s_{|d|}^{|\mathcal{D}|}\}$, where $\mathcal{S}$ contains all the text segments identified in each document $d\in\mathcal{D}$, and every two clusters do not have overlapping text segments (i.e., $E_{i}\cap E_{j}=\emptyset$ for $i\neq j$).
@@
It is important to note that $k$, the number of episodes, is not known in advance and oftentimes, a news article segment may discuss either episodes of a different key event (e.g., an episode with similar aspects that occurred in a different historical key event) or multiple episodes of the current key event. Nonetheless, our goal is to detect the most relevant episodes to the current key event at hand and consequently mine the most distinctive text segments for each of these (hence our constraint of $E_{i}\cap E_{j}=\emptyset$ for $i\neq j$).
4 Methodology